An egg is thrown nearly vertically upward from a point near the cornice of a tall building. It just misses the cornice on the way down and passes a point a distance 43.0 m below its starting point at a time 5.00 s after it leaves the thrower's hand. Air resistance may be ignored.

A) What is the initial speed of the egg?

Express your answer in meters per second.

B)How high does it rise above its starting point?

Express your answer in meters.

C)

What is the magnitude of its velocity at the highest point?

Express your answer in meters per second.

D) What is the magnitude of its acceleration at the highest point?

Express your answer in meters per second squared.

E) What is the direction of its acceleration at the highest point? Up or down.
Answer #1

A) What is the initial speed of the egg?

y - y0 = v0*t - 1/2*g*t^2
with y0 = 0, y = -43 and t = 5.00 s,

-43 = v0 * 5 - 4.9*25

=> v0 = 15.9 m/s^2

B)How high does it rise above its starting point?

v^2 = v0^2 - 2gy

=> 0 = 15.9^2 - 2*9.8y

=> y = 12.9 m

C)What is the magnitude of its velocity at the highest point?

the magnitude of its velocity at the highest point is 0.

D) What is the magnitude of its acceleration at the highest point?

Since the egg falls under gravity, so acceleration due to gravity is constant in magnitude and direction throughout the path of egg.
Hence g = 9.8m/s^2

E) What is the direction of its acceleration at the highest point? Up or down.
Downward